# Joint Bayesian inference of Earth’s magnetic field and core surface flow on millennial timescales

### v1

@@ 1 Introduction
The past two decades of satellite monitoring of the magnetic field, in combination with advancements in numerical simulations of the geodynamo, have generated a wealth of knowledge of processes in Earth’s core on centennial to annual timescales [Gillet2022, Aubert2023, e.g.]. Geodynamo simulations have been able to reproduce the main characteristics of the Earth’s magnetic field and core surface flow over the era of direct observations [Aubert2013, Mound2023, e.g.], e.g. the preference for strong secular variation in the Atlantic hemisphere [Finlay2020]. Comparatively less attention has been concentrated on reproducing characteristic features of geomagnetic field variations over the past $\sim$10,000 years, mostly due to the lack of reference data with adequate resolution and disagreement between models [Wardinski2025]. On these timescales, spanning several convective overturn times ($\sim$130 years), external forcing mechanisms and boundary conditions such as lower mantle heat-flux heterogeneities [Mound2023, e.g.], gravitational coupling between inner-core and mantle and asymmetric inner core growth [Aubert2013, Aubert2017, e.g.], are thought to play a more dominant role.
@@
On millennial timescales, the observational constraints required to investigate core dynamics are provided by indirect records of the geomagnetic field. Through recent technical innovations, models based on these archaeo/palaeomagnetic observations of Earth’s magnetic field are providing information on past field changes with unprecedented resolution [Nilsson2014, Constable2016, Hellio2018, Campuzano2019, Nilsson2022, Schanner2022, e.g.]. Characteristic field observations, including patterns of westward/eastward drift and recurrent field asymmetries, have tentatively been interpreted in terms of large-scale flow [Dumberry2007, Nilsson2020, Nilsson2022, e.g.]. However, due to the many challenges involved, so far only a handful of attempts have been made to actually reconstruct core surface flow based on these field models [Dumberry2006, Wardinski2008, Kianishahvandi2024, Suttie2025, e.g.].
@@
The most recent effort by [Suttie2025], takes advantage of statistics taken from numerical dynamo simulations [Aubert2021] to address key challenges with the contributions of sub-grid errors and magnetic diffusion to induction at the core surface, which can be significant even on short timescales with high-degree field models. The study, which is based on the <pygeodyn> geomagnetic data assimilation package [istas2023, Gillet2019, Huder2019], demonstrates that despite the low spatial and temporal resolution attainable with archaeo/palaeomagnetic data, it is possible to reproduce the large-scale features of modern core surface flow reconstructions over the past century [Gillet2019] as well as changes in length-of-day derived from historical eclipses over the past 1200 years [Morrison2001]. Key ingredients for this are the physical constraints provided by the geodynamo simulations. Although not conclusive, the study also provides evidence that the persistent westward drift of 0.09°/year previously observed in palaeomagnetic field models [Nilsson2020, Nilsson2022] is tied to the long-term average zonal flow in Earth’s core.
@@ 2 Archaeo/palaeomagnetic data
Since the aim is to evaluate a new method to model core surface flow, we will only consider synthetic data in this study. To build this dataset we start from the archaeo/palaeomagnetic data used to construct the pfm9k.2 model [Nilsson2022], including all the pre-treatments implemented for that model. The dataset consists of $N_{ARC}=9440$ archaeomagnetic observations and $N_{SED}=1465$ sedimentary observations from $N_{S}=10$ records covering the past 9000 years. To generate synthetic data we use the Midpath Coupled-Earth dynamo simulation [Aubert2017] and extract a 10,000 year period to represent our reference geomagnetic field and core surface flow over the period 8000 BCE to 2000 CE. The choice of the Midpath simulation is motivated by it’s Earth-like properties, with the ability to resolve geophysical variations on timescales of $\sim$20 years (the Alfvén time in the simulation) and longer, which is enough for the purpose of modelling the core field and flow evolution from archaeo-/palaeomagnetic data. Compared to more recent simulations [Aubert2021, Aubert2023], Midpath has also been run over a comparatively long integration time ($\sim$20,000 years), which will be important for later aspects of the model construction.
Using the reference geomagnetic field, the synthetic dataset was constructed with the identical spatial and temporal distribution as the real data following a similar procedure as described in [Nilsson2021]. All synthetic observations of declination, $D$, inclination, $I$ and intensity, $F$, were generated at the same locations as in the real data but using random (reference) ages consistent with the age uncertainties associated with the respective data. The archaeomagnetic $D,I,F$ observations were then further corrupted with noise, consistent with the given data uncertainties, drawn from a Student’s t distribution with $\nu=4$ degrees of freedom to simulate the presence of outliers. The sediment $D,I$ observations (no sedimentary intensity data were included) were filtered and offset to account for the effects of pDRM, core orientation and inclination shallowing errors, with the associated reference parameters drawn from the respective prior distributions (see section 3.3.2) and then corrupted with Gaussian noise based on the data uncertainties.
@@ 3 Core field and flow model
Figure 1: A schematic chart of the model with data highlighted in blue, model parameters (variables) in red and transformed parameters/quantities in black (for more details see main text). The left-hand part of the chart depicts the iterative construction of the core field and core flow discrete time series, at times $t_{1},...,t_{N_{t}}$. The right-hand part of the chart depicts the likelihood function which is split into two blocks of data: Archeomagnetic data ($k=1,\dots,N_{ARC}$) and Sedimentary data ($k=1,\dots,N_{SED}$, from $s=1,\dots,N_{S}$ records).
Fig. 1 shows a schematic overview of the model with detailed descriptions of each part provided in following three sections.
@@ 3.1 Parameterisation of core field and flow
$\frac{\partial B_{r}}{\partial t}=-\nabla_{h}\cdot(\mathbf{U}_{h}B_{r})+\frac{\eta}{r}\nabla^{2}(rB_{r})$ (1)
with $B_{r}$ the radial magnetic field at the CMB, $\mathbf{U}_{h}$ the core surface flow, $\nabla_{h}\cdot$ the horizontal part of the divergence operator and $\eta$ the magnetic diffusivity. $B_{r}$ and $\mathbf{U}_{h}$ are parameterised in the spectral domain based on spherical harmonic (SH) expansions.
@@
$U_{h}(\theta,\phi)=\nabla\times(T\mathbf{r})+\nabla_{h}(rS)$ (3)
with $T$ and $S$ respectively the toroidal and poloidal scalars. $T$ and $S$ are also expanded in the basis of Schmidt-normalised SH
@@
$\dot{\mathbf{b}}=\textsf{A}(\mathbf{b})\mathbf{u}+\mathbf{e}=\mathbf{f}+\mathbf{e}$ (6)
The first term on the right-hand side (r.h.s.), $\mathbf{f}=\textsf{A}(\mathbf{b})\mathbf{u}$, corresponds to the SV induced by large-scale flow where the matrix A accounts for the Gaunt-Elsasser integrals [Lloyd1990] (as well as the downward continuation of $\mathbf{b}$ to the CMB). The second term on the r.h.s., $\mathbf{e}$, is the error of representativeness, which accounts for the contribution to SV arising from both sub-grid processes and diffusion. The vectors $\mathbf{f}$ and $\mathbf{e}$, of dimension $N_{b}$, contains SH coefficients $\{f_{l}^{m,c},f_{l}^{m,s}\}$ and $\{e_{l}^{m,c},e_{l}^{m,s}\}$ that are defined equivalently to the MF and SV coefficients through equation (2).
Given the inherent limitations of the archaeo/palaeomagnetic data, the large scale (modelled) MF and SV are defined by a truncation at SH degree $L_{b}=5$ [Licht2013, Sanchez2016, Nilsson2022], while the modelled core flow is limited to SH degree $L_{u}=10$ [Suttie2025]. We emphasize that the higher degree flows are retained to ensure that the modelled core flows are consistent with the types of flow seen in the dynamo simulations, but this does not necessarily mean that they can be resolved. The spatial complexity of the flow is further limited by only using the first $N_{v}$ principal components. The principal component analysis is based on a time series of the core surface coefficients from the geodynamo simulation and $N_{v}=46$ (in the case of Midpath) set to retain $95\%$ of the variance. The flow vector can then be written as
$\mathbf{u}=\mathbf{u}_{0}+\Phi\mathbf{v}$ (7)
@@
Contrary to most other models of core flow [Pais2008, Gillet2019, e.g.], ours corresponds to a pure forward model with no need to actually calculate the full $\textsf{A}(\mathbf{b})$ matrix, let alone its inverse form. A considerable speed-up in computation time, which is essential for the HMC implementation, is achieved by instead calculating the whole term $\textsf{A}(\mathbf{b})\mathbf{u}$ directly using products of the magnetic field and the flow in the spatial domain
$\frac{\partial\overline{B_{r}}}{\partial t}=-\overline{B_{r}\nabla_{h}\cdot\mathbf{U}_{h}}-\overline{\mathbf{U}_{h}\cdot\nabla_{h}B_{r}}$ (8)
where the overline denotes the projection onto large length scales. The induced SV is then back-projected onto the spherical harmonics basis. This is similar to how the term $\textsf{A}(\mathbf{b})$ is usually calculated in practice [Lloyd1990], but requires a factor of $2L_{u}$ less calculations. The main benefit, however, is that the design matrices required to go from spectral to spatial domain in (8) for a pre-defined grid, can all be pre-calculated and passed along to the HMC algorithm.
@@ 3.2 Multivariate core field and flow process
We concatenate $\mathbf{v}$ with $\mathbf{\tilde{e}}$ to form an augmented core state vector $\mathbf{z}^{T}=[\mathbf{v}^{T}\mathbf{\tilde{e}}^{T}]$, leaving aside the time-averaged error of representativeness $\mathbf{e}_{0}$ considered known from the dynamo simulation. The temporal evolution of $\mathbf{z}(t)$ and $\mathbf{b}(t)$ are modelled using a time-discrete stochastic process in increments of $\Delta t$ over $N_{t}$ time steps $(t_{1},...,t_{N_{t}})$. The statistical properties of the stochastic process are derived from a prior geodynamo simulation sampled at $\Delta t^{*}=\Delta t=50$ years (in the case of Midpath dynamo series).
@@
Table 1: Prior distributions for model parameters, where $\textsf{I}_{z}$ and $\textsf{I}_{b}$ are identity matrices of rank $N_{z}$ and $N_{b}$ respectively and $\Gamma(\cdot)$ denotes the Gamma distribution.
@@
$\mathbf{z}_{1}=\textsf{L}_{z}\mathbf{w}_{z,1}$ (9)
$\mathbf{b}_{1}=\textsf{L}_{b}\mathbf{w}_{b,1}+\mathbf{b}_{0}$ (10)
where $\textsf{L}_{z}{\textsf{L}_{z}}^{T}=\textsf{K}_{zz}$ and $\textsf{L}_{b}{\textsf{L}_{b}}^{T}=\textsf{K}_{bb}$. In other words, $\textsf{L}_{z}$ and $\textsf{L}_{b}$ are the Cholesky decompositions of $\textsf{K}_{zz}$ and $\textsf{K}_{bb}$. For any given time (in this case $t_{i}=t_{1}$), the core state vector $\mathbf{z}_{i}$ gives access to $\mathbf{u}_{i}$ and $\mathbf{e}_{i}$, which combined with $\mathbf{b}_{i}$ are used to obtain $\dot{\mathbf{b}}_{i}$ through equation (6).
@@
$\mathbf{z}_{i+1}=(\textsf{I}_{z}-\Delta t\textsf{D}_{z})\,\mathbf{z}_{i}+\sqrt{\Delta t}\,\mathbf{r}_{i}$ (11)
@@
$\textsf{K}_{zz^{+}}=$ $E(\mathbf{z}_{i}\mathbf{z}_{i+1}^{T})$ (12)
$\textsf{K}_{\Delta z\Delta z}=$ $E(\Delta\mathbf{z}\Delta\mathbf{z}^{T})$ (13)
@@
$\textsf{D}_{z}=\frac{\textsf{I}_{z}-({\textsf{K}_{zz}}^{-1}\textsf{K}_{zz^{+}})^{T}}{\Delta t}$ (14)
@@
$\textsf{K}_{rr}=\frac{\textsf{K}_{\Delta z\Delta z}}{\Delta t}-\Delta t\textsf{D}_{z}\textsf{K}_{zz}{\textsf{D}_{z}}^{T}$ (15)
@@
Figure 2: Time series and temporal power spectra of (a-f) three core flow coefficients and (g-l) three MF gauss coefficients. The Midpath prior geodynamo series (black) is compared to a 100k long random time series generated with the multivariate core flow and field process (blue), with the shaded blue areas representing the 95% range of 300 random samples. Dashed black lines show the 95% interval ($\pm 2\sigma$) of the prior dynamo series and the dashed red line shows the equivalent range from the ’white noise’ term in equation (24). Black dot-dashed lines corresponds to a spectral index of $p=2$ and $p=4$, for core flow and core field coefficients respectively. Note that the differences in the power spectra towards long periods are due to the limited duration for the Midpath dynamo series.
@@
$E(\mathbf{b^{\prime}}_{i+1}\,|\,\mathbf{x}_{i})=\textsf{K}_{b^{+}x}{\textsf{K}_{xx}}^{-1}\mathbf{x}_{i}\equiv\textsf{D}_{b}\mathbf{x}_{i}$ (16)
@@
$\textsf{K}_{xx}=$ $\begin{bmatrix}\textsf{K}_{bb}&\textsf{K}_{bd}\\ {\textsf{K}_{bd}}^{T}&\textsf{K}_{dd}\end{bmatrix}$ (18)
$\textsf{K}_{b^{+}x}=$ $\begin{bmatrix}\textsf{K}_{b^{+}b}&\textsf{K}_{b^{+}d}\\ \end{bmatrix}$ (19)
@@
$\textsf{K}_{bd}=$ $E\left(\mathbf{b}^{\prime}_{i}\,\dot{\mathbf{b}}_{i}^{T}\right)$ (20)
$\textsf{K}_{dd}=$ $E\left(\dot{\mathbf{b}}_{i}\,\dot{\mathbf{b}}_{i}^{T}\right)$ (21)
$\textsf{K}_{b^{+}b}=$ $E\left(\mathbf{b}^{\prime}_{i+1}\,{\mathbf{b}^{\prime}_{i}}^{T}\right)$ (22)
$\textsf{K}_{b^{+}d}=$ $E\left(\mathbf{b}^{\prime}_{i+1}\,\dot{\mathbf{b}}_{i}^{T}\right).$ (23)
@@
$\mathbf{b}_{i+1}=\mathbf{b}_{0}+\textsf{D}_{b}\mathbf{x}_{i}+L_{bx}\mathbf{w}_{b,i+1}$ (24)
where $\textsf{L}_{bx}{\textsf{L}_{bx}}^{T}=\textsf{K}_{bb|x}$. The second term on the r.h.s. provides the link to $\dot{\mathbf{b}}_{i}$ (through $\mathbf{x}_{i}$) that in turn is related to the core state vector $\mathbf{z}_{i}$ ($\mathbf{u}_{i}$ and $\mathbf{e}_{i}$) through equation (6). Combined with the first term, and for small $\Delta t$, this is effectively equivalent to an Euler scheme, whereas for large $\Delta t$ it converges to the time-average $\mathbf{b}_{0}$. The third ’white-noise’ term on the r.h.s is necessary to account for unmodelled changes in MF and SV and to obtain a stationary distribution for $\mathbf{b}(t)$ (Fig. 2g-l). For large $\Delta t$, the independently parameterised ’white noise’ term will dominate, which means that the recovered variations in $\mathbf{b}(t)$ are partially decoupled from $\mathbf{z}(t)$. However, as demonstrated in Fig. 2, this mostly concerns the higher order and degree Gauss coefficients with short correlation times compared to $\Delta t=50$ years (Fig. 2k), whilst the large-scale and slowly varying magnetic field components (Fig. 2g,j), remain primarily the result of core dynamics captured by $\mathbf{z}(t)$. In other words, the choice of $\Delta t$ determines the degree to which the modelled core surface flow dynamics are linked to the MF variations that we can observe (section 3.3).

### v2

@@ 1 Introduction
The past two decades of satellite monitoring of the magnetic field, in combination with advancements in numerical simulations of the geodynamo, have generated a wealth of knowledge of processes in Earth’s core on centennial to annual timescales [Gillet2022, Aubert2023, e.g.]. Geodynamo simulations have been able to reproduce the main characteristics of the Earth’s magnetic field and core surface flow over the era of direct observations [Aubert2013, Mound2023, e.g.], e.g. the preference for strong secular variation in the Atlantic hemisphere [Finlay2020]. Comparatively less attention has been concentrated on reproducing characteristic features of geomagnetic field variations over the past $\sim$10,000 years, mostly due to the lack of reference data with adequate resolution and disagreement between palaeomagnetic models [Wardinski2025]. On these timescales, spanning several convective overturn times ($\sim$130 years), external forcing mechanisms and boundary conditions such as lower mantle heat-flux heterogeneities [Mound2023, e.g.], gravitational coupling between inner-core and mantle and asymmetric inner core growth [Aubert2013, Aubert2017, e.g.], are thought to play a more dominant role.
@@
On millennial timescales, the observational constraints required to investigate core dynamics are provided by indirect records of the geomagnetic field. Through recent technical innovations, models based on these archaeo/palaeomagnetic observations of Earth’s magnetic field are providing information on past field changes with unprecedented resolution [Nilsson2014, Constable2016, Hellio2018, Campuzano2019, Nilsson2022, Schanner2022, e.g.]. Characteristic field observations, including patterns of westward/eastward drift and recurrent field asymmetries, have tentatively been interpreted in terms of large-scale flow [Dumberry2007, Nilsson2020, Nilsson2022, e.g.]. However, due to the many challenges involved, so far only a handful of attempts have been made to actually reconstruct core surface flow based on these field models [Dumberry2006, Wardinski2008a, Kianishahvandi2024, Suttie2025, Rivera2026, e.g.].
@@
One of the most recent efforts by [Suttie2025], takes advantage of statistics taken from numerical dynamo simulations [Aubert2021] to address key challenges with the contributions of sub-grid errors and magnetic diffusion to induction at the core surface, which can be significant even on short timescales with high-degree field models. The study, which is based on the pygeodyn geomagnetic data assimilation package [Gillet2019, Huder2019, istas2023], demonstrates that despite the low spatial and temporal resolution attainable with archaeo/palaeomagnetic data, it is possible to reproduce the large-scale features of modern core surface flow reconstructions over the past century [Gillet2019] as well as changes in length-of-day derived from historical eclipses over the past 1200 years [Morrison2001]. Key ingredients for this are the physical constraints provided by the geodynamo simulations. Although not conclusive, the study also provides evidence that the persistent westward drift of 0.09°/year previously observed in palaeomagnetic field models [Nilsson2020, Nilsson2022] is tied to the long-term average zonal flow in Earth’s core.
@@ 2 Archaeo/palaeomagnetic data
Since the aim is to evaluate a new method to model core surface flow, we will only consider synthetic data in this study. To build this dataset we start from the archaeo/palaeomagnetic data used to construct the pfm9k.2 model [Nilsson2022], including all the pre-treatments implemented for that model. The dataset consists of $N_{ARC}=9440$ archaeomagnetic observations and $N_{SED}=1465$ sedimentary observations from $N_{S}=10$ records covering the past 9000 years. To generate synthetic data we use the Midpath Coupled-Earth dynamo simulation [Aubert2017] and extract a 10,000 year period to represent our reference geomagnetic field and core surface flow over the period 8000 BCE to 2000 CE. We choose this longer reference time series to be able to generate synthetic data with "true" ages older than 7000 BCE. The choice of the Midpath simulation is motivated by its Earth-like properties, with the ability to resolve geophysical variations on timescales of $\sim$20 years (the Alfvén time in the simulation) and longer, which is enough for the purpose of modelling the core field and flow evolution from archaeo-/palaeomagnetic data. Compared to more recent simulations [Aubert2021, Aubert2023], Midpath has also been run over a comparatively long integration time ($\sim$20,000 years), which will be important for later aspects of the model construction.
Using the reference geomagnetic field, the synthetic dataset was constructed with the identical spatial and temporal distribution as the real data following a similar procedure as described in [Nilsson2021]. All synthetic observations of declination, $D$, inclination, $I$ and intensity, $F$, were generated at the same locations as in the real data but using randomly generated reference ages consistent with the age uncertainties associated with the respective data. The archaeomagnetic $D,I,F$ observations were then further corrupted with noise, consistent with the given data uncertainties, drawn from a Student’s t distribution with $\nu=4$ degrees of freedom to simulate the presence of outliers. The sediment $D,I$ observations (no sedimentary intensity data were included) were filtered and offset to account for the effects of pDRM, core orientation and inclination shallowing errors, with the associated reference parameters drawn from the respective prior distributions (see section 3.3.2) and then corrupted with Gaussian noise based on the data uncertainties.
@@ 3 Core field and flow model
Figure 1: A schematic chart of the model with data highlighted in blue, model parameters (variables) in red and transformed parameters/quantities in black (for more details see main text). The left-hand part of the chart depicts the iterative construction of the core field and core flow discrete time series, at times $t_{1},...,t_{N_{t}}$. The right-hand part of the chart depicts the likelihood function which is split into two blocks of data: Archaeomagnetic data ($k=1,\dots,N_{ARC}$) and Sedimentary data ($k=1,\dots,N_{SED}$, from $s=1,\dots,N_{S}$ records).
Fig. 1 shows a schematic overview of the model with detailed descriptions of each part provided in the following three sections.
@@ 3.1 Parameterisation of core field and flow
$\frac{\partial B_{r}}{\partial t}=-\nabla_{h}\cdot(\mathbf{U}_{h}B_{r})+\frac{\eta}{r}\nabla^{2}(rB_{r}),$ (1)
with $B_{r}$ the radial magnetic field at the CMB, $\mathbf{U}_{h}$ the core surface flow, $\nabla_{h}\cdot$ the horizontal part of the divergence operator and $\eta$ the magnetic diffusivity. The first term on the right-hand side (r.h.s.) represents the advection of the radial field and second term its diffusion. $B_{r}$ and $\mathbf{U}_{h}$ are parameterised in the spectral domain based on spherical harmonic (SH) expansions.
@@
$\mathbf{U}_{h}(\theta,\phi)=\nabla\times(T\mathbf{r})+\nabla_{h}(rS),$ (3)
with $T$ and $S$ respectively the toroidal and poloidal scalars. $T$ and $S$ are also expanded in the basis of Schmidt semi-normalised SH:
@@
$\dot{\mathbf{b}}=\textsf{A}(\mathbf{b})\mathbf{u}+\mathbf{e}=\mathbf{f}+\mathbf{e}.$ (6)
The first term on the r.h.s., $\mathbf{f}=\textsf{A}(\mathbf{b})\mathbf{u}$, corresponds to the SV induced by large-scale flow where the matrix A accounts for the Gaunt and Elsasser integrals [Whaler1986] (as well as the downward continuation of $\mathbf{b}$ to the CMB). The second term on the r.h.s., $\mathbf{e}$, is the error of representativeness, which accounts for the contribution to SV arising from both sub-grid processes and diffusion. The vectors $\mathbf{f}$ and $\mathbf{e}$, of dimension $N_{b}$, contains SH coefficients $\{f_{l}^{m,c},f_{l}^{m,s}\}$ and $\{e_{l}^{m,c},e_{l}^{m,s}\}$ that are defined equivalently to the MF and SV coefficients through eq. (2).
Given the inherent limitations of the archaeo/palaeomagnetic data, the large scale (modelled) MF and SV are defined by a truncation at SH degree $L_{b}=5$ [Licht2013, Sanchez2016, Nilsson2022], while the modelled core flow is limited to SH degree $L_{u}=10$ [Suttie2025]. We emphasize that the higher degree flows are retained to ensure that the modelled core flows are consistent with the types of flow seen in the dynamo simulations, but this does not necessarily mean that they can be resolved. The spatial complexity of the flow is further limited by only using the first $N_{v}$ principal components. The principal component analysis (PCA) is based on a time series of the core surface coefficients from the geodynamo simulation and $N_{v}=46$ (in the case of Midpath) set to retain $95\%$ of the variance. The flow vector can then be written as
$\mathbf{u}=\mathbf{u}_{0}+\Phi\mathbf{v},$ (7)
@@
Contrary to most other models of core flow [Pais2008, Gillet2019, e.g.], ours corresponds to a pure forward model with no need to actually calculate the full $\textsf{A}(\mathbf{b})$ matrix, let alone its inverse form. A considerable speed-up in computation time, which is essential for the HMC implementation, is achieved by instead calculating the whole term $\textsf{A}(\mathbf{b})\mathbf{u}$ directly using products of the magnetic field and the flow in the spatial domain. We use a Gauss-Legendre grid with 16 nodes in $\theta$ (the zeros of the degree-16 Legendre polynomial) and 32 uniformly spaced points in $\phi$. The induced SV is then back-projected onto the spherical harmonics basis using a Gauss-Legendre Quadrature method. This is similar to how the term $\textsf{A}(\mathbf{b})$ is usually calculated in practice [Lloyd1990], but requires a factor of $2L_{u}$ less calculations. The main benefit, however, is that the design matrices required to go from spectral to spatial domain in for a pre-defined grid, can all be pre-calculated and passed along to the HMC algorithm.
@@ 3.2 Multivariate core field and flow process
We concatenate $\mathbf{v}$ with $\mathbf{e^{\prime}}$ to form an augmented core state vector $\mathbf{z}^{T}=[\mathbf{v}^{T}\mathbf{e^{\prime}}^{T}]$, leaving aside the time-averaged error of representativeness $\mathbf{e}_{0}$ considered known from the dynamo simulation. The temporal evolution of $\mathbf{z}(t)$ and $\mathbf{b}(t)$ are modelled using a time-discrete stochastic process in increments of $\Delta t$ over $N_{t}$ time steps $(t_{1},...,t_{N_{t}})$. The statistical properties of the stochastic process are derived from a prior geodynamo simulation sampled at $\Delta t^{*}=\Delta t=50$ years (in the case of Midpath dynamo series).
@@
Table 1: Prior distributions for model parameters, where $\textsf{I}_{z}$ and $\textsf{I}_{b}$ are identity matrices of rank $N_{z}$ and $N_{b}$ respectively and $\Gamma(\cdot)$ denotes the Gamma distribution. See [Nilsson2021] and [Nilsson2022] for justification of the choice of priors for $\nu_{j},\beta_{s},\gamma_{s},\delta D_{s}$ and $\delta I_{s}$.
@@
$\mathbf{z}_{1}=\textsf{L}_{z}\mathbf{w}_{z,1},$ (8)
$\mathbf{b}_{1}=\textsf{L}_{b}\mathbf{w}_{b,1}+\mathbf{b}_{0},$ (9)
where $\textsf{L}_{z}{\textsf{L}_{z}}^{T}=\textsf{K}_{zz}$ and $\textsf{L}_{b}{\textsf{L}_{b}}^{T}=\textsf{K}_{bb}$. In other words, $\textsf{L}_{z}$ and $\textsf{L}_{b}$ are the Cholesky decompositions of $\textsf{K}_{zz}$ and $\textsf{K}_{bb}$. For any given time (in this case $t_{i}=t_{1}$), the core state vector $\mathbf{z}_{i}$ gives access to $\mathbf{u}_{i}$ and $\mathbf{e}_{i}$, which combined with $\mathbf{b}_{i}$ are used to obtain $\dot{\mathbf{b}}_{i}$ through eq. (6).
@@
$\mathbf{z}_{i+1}=(\textsf{I}_{z}-\Delta t\textsf{D}_{z})\,\mathbf{z}_{i}+\sqrt{\Delta t}\,\mathbf{r}_{i},$ (10)
@@
$\textsf{K}_{zz^{+}}=$ $E(\mathbf{z}_{i}\mathbf{z}_{i+1}^{T}),$ (11)
$\textsf{K}_{\Delta z\Delta z}=$ $E(\Delta\mathbf{z}\Delta\mathbf{z}^{T}),$ (12)
@@
$\textsf{D}_{z}=\frac{\textsf{I}_{z}-({\textsf{K}_{zz}}^{-1}\textsf{K}_{zz^{+}})^{T}}{\Delta t}.$ (13)
@@
$\textsf{K}_{rr}=\frac{\textsf{K}_{\Delta z\Delta z}}{\Delta t}-\Delta t\textsf{D}_{z}\textsf{K}_{zz}{\textsf{D}_{z}}^{T}.$ (14)
@@
Figure 2: Time series and temporal power spectra of (a-f) three core flow coefficients and (g-l) three MF gauss coefficients. The Midpath prior geodynamo series (black) is compared to a 100k long random time series generated with the multivariate core flow and field process (blue), with the shaded blue areas representing the 95% range of 300 random samples. Dashed black lines show the 95% interval ($\pm 2\sigma$) of the prior dynamo series and the dashed red line shows the equivalent range from the ’white noise’ term in eq. (23). Black dot-dashed lines corresponds to a spectral index of $p=2$ and $p=4$, (i.e. $P(f)\propto f^{-p}$) for core flow and core field coefficients respectively. Note that the differences in the power spectra towards long periods are due to the limited duration for the Midpath dynamo series.
@@
$E(\mathbf{b^{\prime}}_{i+1}\,|\,\mathbf{x}_{i})=\textsf{K}_{b^{+}x}{\textsf{K}_{xx}}^{-1}\mathbf{x}_{i}\equiv\textsf{D}_{b}\mathbf{x}_{i},$ (15)
@@
$\textsf{K}_{xx}=$ $\begin{bmatrix}\textsf{K}_{bb}&\textsf{K}_{bd}\\ {\textsf{K}_{bd}}^{T}&\textsf{K}_{dd}\end{bmatrix}$, (17)
$\textsf{K}_{b^{+}x}=$ $\begin{bmatrix}\textsf{K}_{b^{+}b}&\textsf{K}_{b^{+}d}\\ \end{bmatrix}$, (18)
@@
$\textsf{K}_{bd}=$ $E\left(\mathbf{b}^{\prime}_{i}\,\dot{\mathbf{b}}_{i}^{T}\right),$ (19)
$\textsf{K}_{dd}=$ $E\left(\dot{\mathbf{b}}_{i}\,\dot{\mathbf{b}}_{i}^{T}\right),$ (20)
$\textsf{K}_{b^{+}b}=$ $E\left(\mathbf{b}^{\prime}_{i+1}\,{\mathbf{b}^{\prime}_{i}}^{T}\right),$ (21)
$\textsf{K}_{b^{+}d}=$ $E\left(\mathbf{b}^{\prime}_{i+1}\,\dot{\mathbf{b}}_{i}^{T}\right),$ (22)
@@
$\mathbf{b}_{i+1}=\mathbf{b}_{0}+\textsf{D}_{b}\mathbf{x}_{i}+L_{bx}\mathbf{w}_{b,i+1},$ (23)
where $\textsf{L}_{bx}{\textsf{L}_{bx}}^{T}=\textsf{K}_{bb|x}$. The second term on the r.h.s. provides the link to $\dot{\mathbf{b}}_{i}$ (through $\mathbf{x}_{i}$) that in turn is related to the core state vector $\mathbf{z}_{i}$ ($\mathbf{u}_{i}$ and $\mathbf{e}_{i}$) through eq. (6). Combined with the first term, and for small $\Delta t$, this is effectively equivalent to an Euler scheme, whereas for large $\Delta t$ it converges to the time-average $\mathbf{b}_{0}$. The third ’white-noise’ term on the r.h.s is necessary to account for unmodelled changes in MF and SV and to obtain a stationary distribution for $\mathbf{b}(t)$ (Fig. 2g-l). For large $\Delta t$, the independently parameterised ’white noise’ term will dominate, which means that the recovered variations in $\mathbf{b}(t)$ are partially decoupled from $\mathbf{z}(t)$. Tests based on the Midpath dynamo series, with various $\Delta t$, shows that the size white noise term is inversely related to MF correlation time $\tau_{b}=\sqrt{\sigma_{b}^{2}/\sigma_{\dot{b}}^{2}}$, where $\sigma_{b}$ and $\sigma_{\dot{b}}$ are the square-root diagonal elements of $K_{bb}$ and $K_{dd}$ respectively. As demonstrated in Fig. 2, the white noise term is mainly important for the higher order and degree Gauss coefficients with short correlation times compared to $\Delta t=50$ years (Fig. 2k), whilst the large-scale and slowly varying magnetic field components (Fig. 2g,j) remain primarily the result of core dynamics captured by $\mathbf{z}(t)$. As a rule of thumb, we find that for $\Delta t=\tau_{b}$, the white noise term added in (23) corresponds to roughly 20% of the total MF variance. In other words, the choice of $\Delta t$ determines the degree to which the modelled core surface flow dynamics are linked to the MF variations that we can observe (section 3.3).